OpenStudy (anonymous):

The deer population in a region 125 000 and the population grows at a rate of 3.2% per year. A) What is the value of b in the exponential relation that models this situation? B) Determine an exponential equation in the form P= ab to the x which models this situation , where P represents the deer population and x represents the number of years. C) Determine the deer population in 15 years? round to the nearest whole number. am soo sorry is my final project and math is one of my lowest and am close to fail the class and not get a credit ):... thanks!!

OpenStudy (anonymous):

A) I'm assuming that b stands for growth rate which would be 1+ percentage of growth/decay = 1+0.032 = 1.032 B) That would be P=125 000 (1.032)^x C) You would want to plug 15 in for x which gives you P= 125 000 (1.032) ^(15) P= 200,495.88 Since you can't have 0.88 of a deer (poor little deer), round to lower answer. In the next 15 years, the population would be 200,495 deers.

OpenStudy (anonymous):

thanks hun!! (: